solve the following recurrence relations

a. 𝑇(𝑛) = 𝑇( 𝑛/4) + 𝑇( 𝑛/2 ) + 𝑛^2

b. T(n) = T(n/5) + T(4n/5) + n

c. 𝑇(𝑛) = 3𝑇( n/4 ) + 𝑐𝑛^2Β 

f. 𝑇(𝑛) = (𝑛/π‘›βˆ’5) * 𝑇(𝑛 βˆ’ 1) + 1

g. 𝑇(𝑛) = 𝑇(log 𝑛) + log 𝑛

h. 𝑇(𝑛) = 𝑇 (𝑛^ 1/ 4) + 1

i. 𝑇(𝑛) = 𝑛 + 7 βˆšπ‘› βˆ™ 𝑇(βˆšπ‘›)

j. 𝑇(𝑛) = 𝑇 ( 3𝑛/4 ) + 1/root(n)


1
Expert's answer
2021-12-20T16:57:05-0500

a.

"T(n)=n^2+5n^2\/16+25n^2\/256+...=n^2\/(1-5\/16)"

"T(n)=O(n^2)"


c.

Comparing with "T(n)=aT(n\/b)+f(n)" :

"a=3,b=4"

so, we have:

"f(n)=\Omega(n^{log_ba})=\Omega(n^{log_43})"

by Master Theorem:

"T(n)=\Theta(f(n))=\Theta(n^2)"


b.


additive term is exactly the current subproblem size, so T(n) is simply the sum of all numbers that appear in this tree. the:

"T(n)\ge nlog_5n" and "T(n)\le nlog_{5/2}n"

so,

"T(n)=\Theta(nlogn)"


g.

for the upper bound:

"T(n)=O(logn)"

also "T(n)=\Omega(logn)"

so,

"T(n)=\Theta(logn)"


j.

using Master Theorem:

"f(n)=O(n^{log_{4\/3}3})"

so,

"T(n)=\Theta(n^{log_{4/3}3})"


h.

"\ud835\udc47(\ud835\udc5b) = \ud835\udc47 (\ud835\udc5b^{ 1\/ 4}) + 1=\ud835\udc47 (\ud835\udc5b^{ 1\/ 16}) + 2=\ud835\udc47 (\ud835\udc5b^{ 1\/ 64}) + 3=\ud835\udc47 (\ud835\udc5b^{ 1\/ 4^n}) + n"

"T(n)=O(n)"


i.

"𝑇(𝑛) = 𝑛 + 7 𝑇(\sqrt𝑛)\sqrt n=7\sqrt n(7T(n^{1/4})n^{1/4}+n^{1/2})+n="


"=7^kn^{1\/2+1\/4+1\/8+...+1\/2^k}T(n^{1\/2^k})+...+n"


"1\/2+1\/4+1\/8+...+1\/2^k=1-1\/2^k"


let "n^{1\/2^k}=2" , then:

"\ud835\udc47(\ud835\udc5b)=7^nT(2)n\/2+...+n"

"\ud835\udc47(\ud835\udc5b)=O(7^nn)"


f.

"\ud835\udc47(\ud835\udc5b) = (\ud835\udc5b\/(\ud835\udc5b\u22125)) \ud835\udc47(\ud835\udc5b \u2212 1) + 1"


"T(n)=(\frac{n}{n-5})^{n-1}T(1)+(\frac{n}{n-5})^{n-2}+...+1"


"T(n)=O(1)"